OpenStudy (anonymous):

(6h)^6 / (6h)^11 divide and simplify

OpenStudy (callisto):

\[\frac{(6h)^6}{(6h)^{11}} = (6h)^{6-11} = \frac{1}{(6h)^{5}} =\frac{1}{(6^5h^5)} \] Can you do the last step? to simplify 6^5

OpenStudy (campbell_st):

the laws for indices when dividing is subtract the powers \[(6h)^{6-11} = (6h)^{-5} \] or \[\frac{1}{(6h)^{5}}\]